### v2

Abstract
The emergence of quantum mechanics and general relativity has transformed our understanding of the natural world significantly. However, integrating these two theories presents immense challenges, and their interplay remains untested. Recent theoretical studies suggest that the single-photon interference covering huge space can effectively probe the interface between quantum mechanics and general relativity. We developed an alternative design using unbalanced Michelson interferometers to address this and validated its feasibility over an 8.4 km free-space channel. Using a high-brightness single-photon source based on quantum dots, we demonstrated single-photon interference along this long-distance baseline. We achieved a phase measurement precision of 16.2 mrad, which satisfied the measurement requirements for a gravitational redshift at the geosynchronous orbit by five times the standard deviation. Our results confirm the feasibility of the single-photon version of the Colella-Overhauser-Werner (COW) experiment for testing the quantum effects in curved spacetime.
The reconciliation of quantum mechanics and general relativity represents a significant challenge in modern physics carlip2001quantum ; will2014confrontation . However, the intersection between these two theories is typically detected under extremely difficult experimental circumstances, such as extremely tiny spatial scales ($10^{-35}$ m) pikovski2012probing , extremely high energy scales ($10^{19}$ GeV) burgess2004quantum , and extremely strong gravitational fields (near black holes) callan1992evanescent . Due to the extreme physical environment required for measurements, indirect tests have been performed, such as those involving Hawking radiation and Penrose superradiation from black holes drori2019observation ; braidotti2022measurement . Nevertheless, thanks to advances in technology and the efforts of theoretical physicists, a series of direct tests have recently been proposed at low-energy scales (within Earth’s gravitational field) using various carriers, including optomechanics pikovski2012probing ; abbott2016observation ; armano2016sub , cold atoms muller2010precision ; pikovski2015universal ; zych2018quantum , atomic and optical clocks hohensee2011equivalence ; delva2018gravitational ; takamoto2020test ; bothwell2022resolving , and photons zych2011quantum ; zych2012general ; rideout2012fundamental ; brodutch2015post ; vallone2016interference ; xu2019satellite ; terno2020large ; mohageg2022deep . Among these various carriers, photons, the fastest information carriers, are weakly coupled to the environment and suitable for large-scale quantum information experiments pan2012multiphoton ; lu2022micius . Tests involving massless particles, such as photons, cannot be interpreted using a Newtonian gravity framework and require a general relativity description zych2012general ; mohageg2022deep . Until now, all experiments that measured the influence of gravity on quantum systems are consistent with non-relativistic Newtonian gravity colella1975observation ; muller2010precision , and all tests of general relativity can be described within the framework of classical (non-quantum) physics pound1960apparent ; shapiro1964fourth ; hafele1972around . The direct experimental test for the evolution of the quantum state under curved spacetime is still lacking. Consequently, the quantum interference of photons serves as a probe to test the interface of quantum mechanics and general relativity zych2011quantum ; zych2012general ; rideout2012fundamental ; brodutch2015post ; vallone2016interference ; terno2020large ; mohageg2022deep .
As a promising avenue for future research, a satellite platform with maneuvering capabilities can potentially explore the gravitational redshift of single photons using single-photon interference rideout2012fundamental ; lu2022micius . For example, Fig. 1(a) illustrates the measurement of single-photon gravitational redshift using a Mach-Zehnder interferometer (MZI). Clearly, on the two paths of the MZI, the blue dashed parts pass through the same gravitational field zych2011quantum ; zych2012general . Meanwhile, the yellow dotted and green dash-dotted portions traverse different gravitational fields, introducing gravitational-induced phase shifts. However, the phase noise introduced by the ground-to-satellite channel in the two arms of the MZI swamps the gravitational redshift phase, rendering the experiment impossible. To overcome this challenge, a Franson interferometer can be utilized in Fig.1(b), consisting of two independent unbalanced Michelson interferometers (UMIs) and a shared free-space channel, equivalent to the MZI rideout2012fundamental . The shared free-space channel enables the cancellation of atmospheric noise as common-mode noise. Assuming a satellite orbiting at an altitude of $h$ and an unbalanced arm of the interferometer of $\delta l$, the redshift phase can be expressed as
$\varphi=\frac{2\pi\delta lg}{\lambda c^{2}}\frac{R_{e}h}{R_{e}+h},$ (1)
where $R_{e}$ is the radius of the Earth, and $\lambda$ is the wavelength of the single photons rideout2012fundamental . The expected observable phase shift in the experiment is 208 mrad, corresponding to a single-photon wavelength of 893.2 nm, with the satellite orbiting at an altitude of 36000 km and the UMI with an unbalanced arm length of 50 m.
Figure 1: (a), In the MZI, single photons coherently travel along the left blue dashed and green dash-dotted paths or the right blue dashed and yellow dotted paths. Since the two blue dashed paths traverse the same gravitational field and cancel each other out, the gravitational redshift effect will introduce the wavelength difference of single photons between the green dash-dotted and yellow dotted paths, resulting in a phase shift in single photon interference zych2011quantum ; zych2012general . (b), To simplify the setup, two blue dashed paths are combined into one, while the yellow dotted and green dash-dotted paths become the unbalanced arm lengths of two UMI in a Franson interferometer rideout2012fundamental .
Recently, Vallone et al. observed interference patterns at the single-photon level, achieving interference visibility of up to 0.67 along the ground-to-satellite channel vallone2016interference . In our current work, we have further concretized the scheme of the satellite-based single-photon version of the COW experiment and have explored its feasibility. To achieve a measurement with at least 5 times the standard deviation (STD) precision, the measurement precision needs to be better than 41.6 mrad. To confirm the feasibility of the experiment and the corresponding technology, it is necessary to perform verification experiments along the horizontal atmospheric channel. Here, we present a demonstration of quantum interference of single photons over an 8.4 km free-space channel, achieving a multi-mode interference visibility of $\mathcal{V}=0.863\pm 0.004$ and a phase measurement accuracy of $\delta\varphi=16.2$ mrad, which meets the accuracy requirement mentioned above. To create a channel representative of the future satellite-ground channel, we chose the horizontal urban atmosphere with an equivalent thickness equal to the vertical atmosphere peng2005experimental . We also developed a high-brightness quantum dot single-photon source (QDSPS) with a brightness exceeding 0.4 GHz and $g^{(2)}(0)=0.071\pm 0.005$. Our demonstration of high-precision phase measurement and high-brightness QDSPS provides a good foundation for future space-based experiments.
In our scheme, single photons are generated from a QDSPS and adjusted to state $\frac{1}{\sqrt{2}}(|H\rangle+|V\rangle)$ using a polarization beam splitter (PBS) and a half waveplate (HWP), as shown in Fig. 2(a). The single photons are then fed into the UMI, which transforms the state to $|\Psi_{t}\rangle=\frac{1}{\sqrt{2}}(|H\rangle|S\rangle+e^{i\varphi_{t}}|V\rangle|L\rangle)$. Here, $|L\rangle$ and $|S\rangle$ denote the states passing through the long and short arms of the UMI, respectively, and $\varphi_{t}$ is a constant internal phase of the UMI at the transmitter. We placed the UMI in a vacuum chamber to achieve high-precision phase measurement and recorded the air pressure values using a computer, which allowed us to isolate the effects of air pressure fluctuations on our measurements. The delay time between $|L\rangle$ and $|S\rangle$ is $\delta t\approx 4$ ns, corresponding to the UMI’s arm difference of $\delta l=c\delta t\approx 1.2$ m ($c$ is the speed of light in vacuum). The single photons were then coupled into single-mode fiber and guided to a launch telescope.
After traveling through the free-space channel in the Shanghai metropolitan area, the single photons were captured by the receiving system, as shown in Fig. 2(b). We used adjustable waveplates to introduce a variable phase $\varphi$ between $|H\rangle$ and $|V\rangle$ state for interference visibility measuring. After passing through the second UMI, the state changes to $|\Psi_{r}\rangle=\frac{1}{\sqrt{2}}|L\rangle|S\rangle(|H\rangle+e^{i(\varphi_{t}+\varphi+\varphi_{r})}|V\rangle)$, where $\varphi_{r}$ is the phase induced by the UMI at the receiver. Finally, we measured the photon under the Pauli operator $\sigma_{x}$ using an HWP and a PBS, collected photons with multi-mode fibers, with a mode field diameter of 105 µm, supporting up to approximately 40000 different guided modes, and detected using single-photon avalanche detectors (SPAD).
Figure 2: Illustration of experimental setup. (a), The QDSPS and UMI generate single photons in a time-bin qubit, which is crucial for achieving high-precision measurements. b, The optical receiving system is mounted on a high-load pitch and yaw platform for coarse pointing. The telescope has an aperture of 127 mm and a beam reduction ratio of 37.5. The received 671-nm-wavelength laser is separated by a dichroic mirror (DM) and imaged by a camera for fine tracking of the fast steering mirror. A motorized rotation stage controls the three waveplates to achieve dynamic modulation of the interference phase. We also use a 10-nm-wide interference filter (IF) to reduce the dark count. (c), The scheme and photo of the UMI. All optical elements adhere to an optical substrate that measures 160 mm in length, 80 mm in width, and 30 mm in height. To keep the UMI compact, we adopt a round-trip optical design and append imaging systems in the long arm of the UMI. The red arrow represents the direction of the beam propagation in a photo of the UMI, and the interferometer is put in a vacuum chamber with a temperature controlling system. QWP, quarter waveplate; mFPC, manual fiber polarization controller; FSM, fast steering mirror.
A high-brightness, high-purity single-photon source is crucial for mitigating the significant channel attenuation that occurs in long-distance free-space communication channels. Compared to heralded single-photon sources, which generate photons probabilistically through spontaneous parametric down-conversion, solid-state single quantum emitters can produce high-quality single photons with greater efficiency and reliability. Specifically, we employ a self-assembled In(Ga)As quantum dot embedded in a 2.5 µm micropillar as a near-perfect single-photon source he2013demand ; ding2016demand ; deng2019quantum . In our experiment, we excite the quantum dot with a narrow-linewidth continuous-wave laser, which drives the quantum dot resonantly. We collect the resonance fluorescence through a background-free confocal setup with an extinction ratio of about $110^{7}:$ 1. The count rate of single photons at 893.2 nm wavelength is more than 0.4 GHz, which reduces to about 10 kHz due to complete system attenuation. The purity of single photons is characterized by the second-order correlation, where the antibunching signature shows a raw $g^{(2)}(0)=0.071\pm 0.005$, indicating the inherent property of resonance fluorescence. Further details are provided in the Methods.
The ultra-stable field-widened UMI serves as the fundamental basis for this experiment. To ensure the compactness and stability of the UMI’s optical path difference, we utilized ultraviolet-curing optical adhesives to integrate all optical elements onto an optical bench. These adhesives were chosen due to their low shrinkage and stress, which are important factors in preserving the stability of the UMI. Research has shown that fused silica, which has a low thermal expansion coefficient, is the optimal material for all optical elements used in this experiment li2022quantum .
Due to the unbalanced arm length in UMI, its interference performance is significantly affected by atmospheric turbulence, including incident angle jitter and wavefront distortion. This effect can cause reduced interference visibility and degraded phase stability. The spatial filter, such as single-mode fiber coupling, is generally effective, while it entails much more significant attenuation. An alternative solution is to incorporate an imaging system into the UMI. This approach can address the above issues while maintaining good multi-mode interference performance jin2018demonstration ; jin2019genuine . The imaging system comprises two lenses and reduces mode-dependent path length, as illustrated in Fig. 2(c). Thus, it ensures high interference visibility and phase stability despite temporal and spatial photon distortion between different arms. Since photons in the long arm pass through the imaging system twice, any spatial mode transformation becomes a nearly identical matrix that similarly affects photons in both arms. To evaluate the efficacy of this imaging system, we introduced an angular change of 1.6 mrad to the incident light of the UMI. Consequently, there was a variation of 0.04 mrad and 0.06 mm in the overlap of the beam’s direction and position, respectively, for the beam passing through the long and short arms, which are much smaller than the spot diameter of 2 mm and the divergence angle of 0.55 mrad.
This experiment aims to achieve highly accurate phase measurements by carefully isolating the interference phase from various sources of noise, as summarized in Table 1. Our investigation identifies six principal sources of noise, namely the photon’s center wavelength, air pressure, temperature, atmospheric turbulence, shot noise, and inconsistency of SPADs. Table 1 reveals that temperature is the predominant noise source affecting long-term stability, while shot noise and inconsistency of SPADs impact short-term stability. Shot noise arises from the Poisson statistics of photons and can only be mitigated by increasing the number of received photons. The inconsistency of SPADs primarily introduces phase noise under different channel attenuations. Furthermore, temperature and air pressure emerge as key environmental noise sources, as they introduce thermal shrinkage of interferometers and instabilities in the optical path due to variations in air pressure ciddor1996refractive . To address these challenges, we placed two unbalanced Michelson interferometers (UMIs) in separate environments and implemented passive heat insulation and chamber temperature control. Despite efforts to regulate temperature, imprecise control led to long-term drift in the interference phase of the UMI. Additionally, to mitigate phase noise caused by air pressure fluctuations, we maintained a vacuum environment with pressures below 0.5 Pa. We also categorized the effect of atmospheric turbulence on phase noise into transverse and axial components, with an integrated imaging system in UMI effectively suppressing the transverse component. Simulation results indicated that axial turbulence does not significantly contribute to the overall noise. Furthermore, our findings highlight that the drift in the photon’s center wavelength is not the primary source of noise. These insights shed light on the intricate interplay between noise factors and enable improved strategies for achieving high precision in interferometric phase measurements. Further details on the experimental analysis are provided in the Methods.
Table 1: Total noise analysis.
Noise source Corresponding phase noise
Photon’s center wavelength 0.002 mrad/day
Air pressure (transmitter) 0.08 mrad
Air pressure (receiver) 0.1 mrad
Temperature 0.137 mrad/s/K
Atmospheric turbulence (transverse) 0.3 mrad
Atmospheric turbulence (axial) 0.001 mrad
Shot noise 4.3 mrad
Inconsistency of SPADs 15.6 mrad
[…]
Figure 3: Experimental results. The sampling rate is set as 0.1 Hz for accumulating enough counts to weaken shot noise. a, Experimental measurement of multi-mode interference visibility along free-space channel by scanning phase $\varphi$ with 2 cycles. The scatter spots indicate the measured value, whose bars denote shot noise. The dashed line indicates the sine fitting curve, which shows the interference visibility $\mathcal{V}=0.863\pm 0.004$. b, Experimental measurement of phase measurement precision along free-space channel with fixed phase $\varphi$. The red scatter spots indicate the measured value, whose bars denote shot noise. The blue dashed line indicates the linear fitting result, which shows the long-term stability of phase measurement with slope $0.117\pm 0.006$ mrad/s. The shaded area shows the the STD of phase measurement without slope as short-term stability.
We measured the multi-mode interference visibility of single-photon interference in an 8.4-km horizontal atmosphere channel located in Shanghai. Here, we use $\frac{C_{1}-C_{2}}{C_{1}+C_{2}}$ to measure interference visibility, where $C_{1}$ and $C_{2}$ represent the single-photon counts of two SPADs. By adjusting $\varphi$, $C_{1}$ can be optimized to reach its maximum value, at which point $\frac{C_{1}-C_{2}}{C_{1}+C_{2}}$ becomes the interference visibility $\mathcal{V}$. In Fig. 3(a), three wave plates in the receiving system introduce dynamic phase $\varphi$ for multi-mode interference. The measured multi-mode interference visibility $\mathcal{V}=0.863\pm 0.004$, indicates the effectiveness of the imaging system against atmospheric turbulence. By directly measuring the phase at a frequency of 0.1 Hz under fixed phase $\varphi$, we achieved a phase measurement precision of 35.8 mrad in Fig. 3(b). However, a long-term linear phase drift was observed and found to be $0.117\pm 0.006$ mrad/s, which is shown as a blue dashed line. Our previous noise analysis revealed that the temperature drift of the UMI causes long-term phase drift, calculated as 0.137 mrad/s/K. To maintain temperature stability, the air conditioner was set to a constant nighttime temperature during the experiments. Without the long-term phase drift, the short-term stability of phase measurement precision 16.2 mrad is shown as a gray area. Through noise analysis, we identified the inconsistency of SPADs and shot noise as the primary sources of noise. The shot-noise-induced phase noise was calculated to be 4.3 mrad. The inconsistency of SPADs would inevitably introduce phase noise under various free-space channel attenuation, estimated as 15.6 mrad. More details are provided in the Methods.
For the satellite-based experiment, it is crucial to mitigate all noise sources effectively. One type of noise, the long-term phase drift caused by temperature stability, can be mitigated through several approaches, such as more sophisticated active temperature control and passive thermal insulation using multiple layers armano2019temperature . Additionally, employing ultra-low-expansion (ULE) glass as the optical bench for the UMI, which has a coefficient of thermal expansion two orders of magnitude lower than that of the fused silica at the specific temperature, can further mitigate the impact of temperature drift on phase stability. Secondly, the inconsistency of SPADs and shot noise affects short-term stability. To mitigate this issue, we will adopt a time-division phase measurement approach with a single SPAD. Notably, for the satellite-based experiment, several issues warrant consideration, such as Doppler noise caused by satellite motion and so on. Furthermore, calibration for $\varphi_{r}$ and $\varphi_{t}$ are imperative for both UMIs located on the satellite and at the ground station. More details are provided in the Methods.
Generally, we developed high-brightness high-purity QDSPS and high-precision single-photon phase measurement technology for future satellite-based experiments. Therefore, we believe that our current work combined with the pioneering work of Vallone et al. vallone2016interference have collectively demonstrated the feasibility of employing single-photon interference to measure gravitational redshift, thereby testing the interface of general relativity and quantum mechanics. Moreover, our results identify the primary noise sources to facilitate the future use of corresponding methods to reduce noise and improve measurement precision.
[…]
Acknowledgments We acknowledge insightful discussions with H. Wang, J. Qin, T. Zeng and L. Huang. This work has been supported by the National Key R&D Program of China (Grants No. 2020YFA0309803, 2020YFC2200103), the National Natural Science Foundation of China (Grants No. 11904358, 12174374, and 12274398), the Innovation Program for Quantum Science and Technology (Grants No. 2021ZD0300100), the Chinese Academy of Sciences (CAS), Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01), the Natural Science Foundation of Shanghai (Grants No. 20ZR1473700), the Shanghai Rising-Star Program (Grants No. 21QA1409600) and Anhui Initiative in Quantum Information Technologies, the CAS Project for Young Scientists in Basic Research (Grant No. YSBR-046). Y. Cao and Y.-H. Li were supported by the Youth Innovation Promotion Association of CAS (under Grant No. 2018492, 2023475).